OpenStudy (angel_kitty12):

Solve for x and y. 3x−y+19i=1+(2x+3y)i

OpenStudy (angel_kitty12):

I'm confused on where to start.

OpenStudy (angel_kitty12):

Imaginary number

OpenStudy (steve816):

ohh ok, than that makes stuff complicated.

OpenStudy (angel_kitty12):

Exactly

OpenStudy (agent0smith):

The real parts, and the imaginary parts, on each side need to be equal. The real part on the left is 3x-y. It needs to be equal to the real part on the right, which is 1. The imaginary part on the left is 19. It needs to be equal to the imaginary part on the right, which is 2x+3y.

OpenStudy (angel_kitty12):

Okay I understand that

OpenStudy (angel_kitty12):

3x-y=1 (2x+3y)i=19i

OpenStudy (angel_kitty12):

Now what? Do I plug in y?

OpenStudy (angel_kitty12):

Y=3x-1

OpenStudy (angel_kitty12):

19-2x=3y

OpenStudy (angel_kitty12):

19-2x=3(3x-1)

OpenStudy (angel_kitty12):

?

OpenStudy (agent0smith):

Yes, solve for x. Then you can use it to find y.

OpenStudy (angel_kitty12):

X=2

OpenStudy (angel_kitty12):

Y=5

OpenStudy (angel_kitty12):

Thank you so much